OpenStudy (anonymous):

Find the dimensions of a rectangle whose perimeter is 16 meters and whose area is 15 square meters.

OpenStudy (anonymous):

Ok solve the system 2a + 2b =16; a*b =15

OpenStudy (anonymous):

you have 2 equations and two unknowns, a and b. Have at it!!

OpenStudy (anonymous):

So far on paper I have x times x+16 = 15 x (x+16) = 15 x^2 + 16x = 15 x^2 + 16x - 15 = 0 (x+ ) (x - ) = 0 shall I put (x + 5) (x - 3) = 0

OpenStudy (dan815):